### v2

# Towards Full-scene Domain Generalization in Multi-agent Collaborative Bird’s Eye View Segmentation for Connected and Autonomous Driving

Abstract
Collaborative perception has recently gained significant attention in autonomous driving, improving perception quality by enabling the exchange of additional information among vehicles. However, deploying collaborative perception systems can lead to domain shifts due to diverse environmental conditions and data heterogeneity among connected and autonomous vehicles (CAVs). To address these challenges, we propose a unified domain generalization framework applicable in both training and inference stages of collaborative perception. In the training phase, we introduce an Amplitude Augmentation (AmpAug) method to augment low-frequency image variations, broadening the model’s ability to learn across various domains. We also employ a meta-consistency training scheme to simulate domain shifts, optimizing the model with a carefully designed consistency loss to encourage domain-invariant representations. In the inference phase, we introduce an intra-system domain alignment mechanism to reduce or potentially eliminate the domain discrepancy among CAVs prior to inference. Comprehensive experiments substantiate the effectiveness of our method in comparison with the existing state-of-the-art works. Code will be released at https://github.com/DG-CAVs/DG-CoPerception.git.
Index Terms:
Domain generalization, vehicle-to-vehicle collaborative perception, autonomous driving, bird’s eye view segmentation.
I Introduction
Recently, multi-agent collaborative perception liuWhen2comMultiAgentPerception2020 ; huWhere2commCommunicationEfficientCollaborative2022 ; yangWhat2commCommunicationefficientCollaborative2023 ; liuWho2comCollaborativePerception2020 ; xuV2XViTVehicletoEverythingCooperative2022 ; chenVehicleServiceVaaS2023 has attracted increasing attention in the autonomous driving community, due to its promising way to overcome the limitations of a single-agent perception system. For example, a single-agent perception system may suffer from occlusion and sparse sensor observation at a far distance, which may lead to a high risk of traffic accidents. Collaborative perception can address these issues by employing vehicle-to-vehicle (V2V) collaboration, where visual information (e.g., raw sensory information, intermediate perception features, and perception results) from multiple nearby CAVs can be shared to improve the accurate understanding of the environment.
In collaborative perception, current approaches mainly aim to strike a balance between performance and bandwidth consumption while developing new perception architectures. For example, Liu et al. liuWho2comCollaborativePerception2020 leveraged a three-step handshake communication protocol to determine the information of which CAVs should be shared. Chen et al. chenCooperCooperativePerception2019 conducted a study to implement collaborative perception by fusing raw LiDAR point clouds of different CAVs. Xu et al. xuV2XViTVehicletoEverythingCooperative2022 proposed a new vision transformer framework which can fuse information of the on-road vehicles and the roadside infrastructure. Hu et al. senkangAdaptiveCommunicationsCollaborative2023 proposed an adaptive communication scheme which can construct the communication graph and minimize the communication delay according to different channel information states.
While the previously mentioned methods have made significant progresses in balancing performance and bandwidth while optimizing communication graph construction, they regrettably overlook a critical challenge - the domain shift problem in collaborative perception. These methods typically assume that both training and testing data originate from the same domain, a condition rarely met in real-world applications. The consequence of neglecting the domain shift is a severe deterioration in perception performance when these methods are implemented in environments characterized by domain variations. For instance, in OPV2V xuOPV2VOpenBenchmark2022 and nuScenes caesarNuScenesMultimodalDataset2020 , the widely recognized benchmarks for bird’s eye view (BEV) semantic segmentation, their data collection settings exhibit notable disparities, encompassing differences in intrinsic and extrinsic parameters. One dataset is sourced from a simulator, while the other derives from real-world observations. When a model initially trained on OPV2V is directly transferred to the nuScenes dataset, a substantial performance drop is observed. Consequently, there is an urgent need to achieve domain generalization for collaborative perception.
Domain generalization (DG) is an active research topic with numerous methods liDomainAdaptationBased2023 ; zhouDomainGeneralizationMixStyle2021 ; liDomainGeneralizationAdversarial2018 ; yangFDAFourierDomain2020 ; volpiGeneralizingUnseenDomains2018 . However, when applied to collaborative perception in the context of CAVs, the domain generalization problem becomes notably more intricate. Several challenges make collaborative perception in CAVs a demanding domain for DG. First, for collaborative BEV segmentation, there are few open-source benchmarks, and annotating data is resource-intensive and time-consuming. Only OPV2V xuOPV2VOpenBenchmark2022 and V2X-Sim liV2XSimMultiAgentCollaborative2022 provide accessible datasets. Moreover, CAV environment is highly dynamic, subject to frequent changes. For example, domains shift when transitioning from city roads to countryside, and similar shifts are also observed when moving from daytime to nighttime conditions. Furthermore, data heterogeneity among collaborative vehicles also presents a significant challenge. In practical scenarios, different vehicles may operate under diverse environmental conditions during collaborative perception. For instance, one vehicle might navigate in darkness, while another is positioned in a well-lit area. The absence of standardized parameters across different onboard cameras lead to variations in image quality, affecting factors like brightness, contrast, and color zhouDomainGeneralizationSurvey2022 . As a result, these divergent environmental conditions and sensor characteristics inevitably create a substantial domain gap between vehicles, leading to performance degradation in collaborative perception.
To tackle these challenges, we present a comprehensive domain generalization framework that operates seamlessly in both the training and inference phases of collaborative perception when machine learning is used. Firstly, our approach starts with the development of an amplitude augmentation method, a resource-efficient technique that does not need extra annotated data and is ready for immediate use. Next, we introduce a meta-consistency training scheme rooted in meta-learning principle. This scheme equips the model with the ability to generalize effectively to unseen domains while fostering the acquisition of domain-invariant features. Finally, our framework includes an intra-system domain alignment scheme, strategically designed to mitigate data heterogeneity and bridge the domain gaps among various CAVs. The main contributions of this paper are summarized as follows.
  • We address the practical problem of domain generalization for collaborative perception. To the best of our knowledge, this is the first work to study domain generalization in the area of collaborative perception.
  • We delve into the underlying reasons for domain shifts at the image level and introduce an amplitude augmentation method to enhance the low-level distribution of images. Additionally, we collect a dataset that aids in augmenting the source domain for application to different target domains.
  • We establish a meta-consistency training scheme based on meta-learning which guides the model to adapt to new domains. We also carefully design a consistency loss to constrain the feature distribution between the features during meta-training and meta-testing phases, which can encourage the model to learn the domain-invariant representations.
  • We devise an intra-system domain alignment scheme tailored to reduce data heterogeneity and bridge domain gaps among diverse CAVs.
The remainder of this paper is organized as follows. In Sec. II, we review the related work of collaborative perception and domain generalization. In Sec. III, we introduce our proposed domain generalization method in detail. In Sec. IV, we conduct comprehensive experiments to evaluate our method. Finally, we conclude our work in Sec. V.
II Related Work
II-A Collaborative Perception
Despite the notable progress in autonomous driving in recent years, single-agent perception systems face significant challenges related to occlusions and limitations of sensor coverage. For example, in a crowded urban environment, a vehicle may be blocked by other vehicles, pedestrians, or buildings, which makes it difficult to perceive its surrounding environment. Multi-agent collaborative perception emerged as a solution to addressing these issues wangV2VNetVehicletoVehicleCommunication2020 ; liuWho2comCollaborativePerception2020 ; liuWhen2comMultiAgentPerception2020 ; huWhere2commCommunicationEfficientCollaborative2022 . Based on data sharing strategies, collaborative perception can generally be divided into three aspects: 1) early fusion chenCooperCooperativePerception2019 where raw data is shared with ego vehicle; 2) intermediate fusion where intermediate features extracted from each vehicle’s sensor data are transmitted and shared; and 3) late fusion where perception outputs are shared (e.g., bounding box position, segmentation mask, etc.). Recent works indicate that intermediate fusion is more efficient than early fusion and late fusion because it effectively balances accuracy and communication bandwidth. For example, Liu et al. liuWho2comCollaborativePerception2020 proposed a handshake communication mechanism, which consists of three stages, namely request, match, and connect, which can determine which agents should communicate. Li et al. liLearningDistilledCollaboration2021 leveraged knowledge distillation to train a model by a teacher-student framework and proposed a kind of matrix that allows the agent to adaptively highlight the information region, thereby achieving effective intermediate fusion. Wang et al. wangV2VNetVehicletoVehicleCommunication2020 proposed an intermediate fusion strategy where all agents transmit features derived from the raw point cloud to strike the balance between bandwidth and precision. Despite the aforementioned advances of fusion methods, the domain shift caused by different environmental conditions and data heterogeneity among CAVs has not been investigated well in collaborative perception.
II-B Domain Generalization
The domain shift problem has seriously hindered large-scale deployments of machine learning models. To tackle this issue, a lot of DG methods have been proposed gongDLOWDomainFlow2019 ; ghifaryDomainGeneralizationObject2015 ; liEpisodicTrainingDomain2019 ; yangFDAFourierDomain2020 ; liFeatureCriticNetworksHeterogeneous2019 . DG is dedicated to learning a model trained on multiple source domains that can generalize to unseen target domains. For example, Motiian et al. motiianUnifiedDeepSupervised2017 leveraged a contrastive loss to learn an embedding subspace that is discriminative, and the mapped visual domains are semantically aligned and yet maximally separated, which requires only a few labeled target samples. Some other domain generalization methods are based on data augmentation, which is a common practice to regularize the training of machine learning models to avoid overfitting and improve generalization, e.g., Volpi et al. volpiAddressingModelVulnerability2019 defined a new data augmentation rule to transform images with a different distribution where the model is most vulnerable, thereby making the model more robust against distribution shifts. Volpi et al. volpiGeneralizingUnseenDomains2018 also proposed a data augmentation method to leverage adversarial gradients obtained from a task classifier to perturb input images. In addition, meta-learning, also known as learn-to-learn, is another method helpful for domain generalization, which exposes a model to domain shift and guides the model to learn how to learn from the domain shift. Finn et al. finnModelAgnosticMetaLearningFast2017 proposed MAML, a method that partitions the training data into meta-train and meta-test subsets, and trains a model using the meta-train subset to enhance its performance on the meta-test subset. Our method also utilizes meta-learning scheme as our foundational learning scheme. However, we propose a novel amplitude augmentation method to simulate the domain shift, and leverage a consistency optimization to enable the model to learn domain-invariant features.
Figure 1: Overall Architecture. 1) The first part of our proposed method is amplitude augmentation which can transform the original source domain data to different target domains. 2) The second part is meta-consistency training, which can simulate the domain shift, guiding the model to learn how to learn from different domains. Then, we exploit the meta-consistency loss to encourage our model to learn the domain-invariant features, thereby enhancing the ability of generalization of the model. 3) The third part is intra-system domain alignment, which can minimize the domain gap among the data perceived by different collaborative vehicles prior to inference.
III Our Method
III-A Problem Formulation and Overview
Collaborative Domain Generalization. Let $(\mathcal{X},\mathcal{Y})$ denote the input and label space in a collaborative BEV segmentation task, $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{K}\}$ be the set of $K$ source domains. In each domain, $\mathcal{D}_{k}=\{(x_{i}^{(k)},y_{i}^{(k)})\}^{N_{k}}_{n=1}$ is a set of samples from domain distribution ($\mathcal{X}_{k},\mathcal{Y}$), where $x_{i}^{(k)}$ is the input image set of CAVs at a certain time, $y_{i}^{(k)}$ is the corresponding label, $N_{k}$ is the number of samples in domain $k$. The goal is to learn a collaborative BEV segmentation model $F:\mathcal{X}\rightarrow\mathcal{Y}$ based on the data from multiple source domains $\mathcal{D}$, which can be generalized to the unseen target domain.
Overview. Our approach offers a unified domain generalization framework for both training and inference in collaborative perception, making it broadly applicable to encoder-decoder structures. Specifically, we first propose an amplitude augmentation (AmpAug) method, which enhances the model’s robustness against domain discrepancies in CAV driving. AmpAug leverages fast Fourier transform to convert images into the frequency domain, incorporating low-frequency signals from our target dataset’s amplitude spectrum. This process culminates in a new synthetic amplitude spectrum, seamlessly integrated with the source domain’s phase spectrum. The result is an augmented image achieved through inverse Fourier transform.
Following AmpAug, we propose our meta-consistency training scheme, a method capable of simulating domain shifts and optimizing the model using a carefully crafted consistency loss function. This loss drives the encoder to learn domain-invariant features by minimizing the maximum mean discrepancy of the features obtained during both meta-training and meta-testing phases. Finally, our intra-system domain alignment mechanism is introduced, aimed at mitigating, or even eliminating, domain discrepancies among connected and autonomous vehicles (CAVs) before the inference stage. This is achieved by translating image styles and unifying the distribution of image pixels in color space. Each of these components will be discussed in detail in the following sections. The overall framework is shown in Fig. 1.
III-B Amplitude Augmentation
Figure 2: Visualization of our constructed target dataset $\mathcal{D}_{t}$.
As for collaborative perception, the environments sensed by CAVs are extraordinarily complex due to dynamic transformations across time and space. Factors such as varying weather conditions, diverse driving scenarios, fluctuating illuminations, and distinct image colors contribute to this complexity. These environmental fluctuations result in domain discrepancy, significantly impeding perception performance when deploying models in real-world settings.
Our research, upon careful examination, has revealed that this domain discrepancy primarily arises from variations in the low-frequency spectrum of images. Importantly, these variations do not affect the image’s finer details, encompassing high-frequency spectrum elements like objects, cars, and people. Furthermore, these low-frequency spectrum variations hold substantial sway over the image. If these variances are not adequately represented in the training dataset, the models fail to generalize effectively. Consequently, we introduce a straightforward yet highly effective method known as AmpAug. AmpAug serves the purpose of augmenting these low-frequency variations, compelling the model to acquire domain-invariant representations. This augmentation, in turn, reduces the model’s susceptibility to the influence of low-frequency spectrum variations in images.
First, we construct a dataset $\mathcal{D}_{t}$, which is carefully collected under different weather conditions (including sunny, rainy, foggy days), driving scenarios, colors, etc. The visualization of our constructed target dataset $\mathcal{D}_{t}$ is shown in Fig. 2. The reason for constructing this dataset is that we want to obtain enough different amplitude spectrum of the images and leverage it to augment the amplitude spectrum of the images in the source domain dataset $\mathcal{D}_{s}$.
Specifically, consider our constructed dataset $\mathcal{D}_{t}=\{x_{i}^{t}\}^{N_{t}}_{i=1}$ and a source domain dataset $\mathcal{D}_{s}=\{x_{i}^{s},y_{i}^{s}\}^{N_{s}}_{i=1}$, where $x_{i}^{t},x_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ ($H$ and $W$ are the height and width of the image, respectively; $C$ is the number of channels), $y_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ is the corresponding label. We first sample an image $x_{i}^{t}$ from our constructed dataset $\mathcal{D}_{t}$, and convert it to the frequency space by fast Fourier transform (FFT):
$\mathbf{F}(x_{i}^{t})(u,v,c)=\sum^{H-1}_{j=0}\sum^{W-1}_{k=0}x_{i}^{t}(h,w,c)e^{-2j\pi\left(\frac{h}{H}u+\frac{w}{W}v\right)}$
After obtaining the frequency representation $\mathbf{F}(x_{i}^{t})$, we can decompose it into amplitude spectrum $A(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$ and phase spectrum $P(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$, which represents the low-level distributions (e.g., style, light, etc.) and high-level distributions (e.g., details and object, etc.), respectively yangFDAFourierDomain2020 . Then, we sample another image $x_{j}^{s}$ from the source dataset and perform the same operation to obtain the amplitude spectrum $A(x_{j}^{s})$ and phase spectrum $P(x_{j}^{s})$. After that, we can obtain the augmented image $x_{j}^{s\rightarrow t}$ by incorporating the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$.
We first introduce a binary mask $M$ which can be used to control the proportion of the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$ in the augmented image $x_{j}^{s\rightarrow t}$, $M=(h,w)=\mathbf{1}_{h\times w}$ where $h\in[-\alpha H:\alpha H]$, $w\in[-\alpha W:\alpha W]$, $\alpha\in(0,1)$. Then, we generate a new amplitude spectrum distribution by as follows:
$A(x_{j}^{s\rightarrow t})=(I-M)\cdot A(x^{s}_{j})+M\cdot A(x^{t}_{i})$ (1)
where $x^{s}_{j}$ and $x^{t}_{i}$ are sampled from source domain dataset $\mathcal{D}_{s}$ and target domain dataset $\mathcal{D}_{t}$, while $I$ denotes the identity matrix. After obtaining the synthetic amplitude spectrum, we integrate it with the source domain phase spectrum to generate the augmented image by inverse Fourier transform $\mathbf{F}^{-1}$:
$x^{s\rightarrow t}_{i}=\mathbf{F}^{-1}\left(A(x_{j}^{s\rightarrow t}),P(x^{s}_{i})\right)$ (2)
Discussion. Here we elaborate the rationale behind the proposed method. Several factors contribute to this choice, with the first being its independence of annotation. In those scenarios related to CAVs, acquiring labeled data is both costly and labor-intensive, often cited as a bottleneck in research and development, and posing challenges for both industry and academia, especially when accounting for time-varying weather conditions (e.g., sunny, rainy, foggy, snowy), color discrepancies, and diverse driving scenarios. Within the domain of collaborative BEV segmentation, open source datasets are scarce. To our best knowledge, only two datasets exist xuOPV2VOpenBenchmark2022 ; liV2XSimMultiAgentCollaborative2022 , both of which are derived from simulations rather than real-world environments. The lack of real-world data underscores the urge need for developing methods that can operate effectively in the absence of extensive annotated datasets. Based on this observation and restriction, we propose AmpAug without using additional annotation. The second notable advantage of our proposed method is its plug-and-use nature. This feature is particularly advantageous as it allows for effortless integration into existing training pipelines without any modifications. This plug-and-use characteristic not only enhances the usability of the method but also significantly reduces the time and resources required for implementation.
III-C Meta-consistency Training
Meta training. We use the gradient-based meta-learning algorithm as our foundational learning scheme, which can learn generalizable model by simulating the real-world domain shift in the training phase. In our settings, the domain shift comes from the data generated from the frequency space with amplitude augmentation. Specifically, in each iteration, we consider the raw camera data perceived by CAVs’ cameras as meta training data, and its counterparts generated from frequency space as meta testing data. The meta-learning can be divided into two stages. We first update the model parameters $\theta$ using the meta-train data $\mathcal{D}_{s}$ with cross entropy $\mathcal{L}_{CE}$:
$\mathcal{L}_{CE}=\frac{1}{N}\sum^{N-1}_{i=0}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right)$ (3)
where $y_{i}$ is the ground truth label and $\hat{y}_{i}$ is the predicted segmentation map. The model parameters $\theta$ are updated by the gradient descent:
$\theta^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{CE}(\mathcal{D}_{s};\theta)$ (4)
where $\alpha$ is the learning rate in the inner loop update. Then, we leverage a meta-learning step to enhance the generalization ability of the model by simulating the domain shift in the meta-test data $\mathcal{D}_{t}$. Specifically, in the meta-learning phase, the meta-objective $\mathcal{L}_{meta}$ is computed with the updated parameters $\theta^{\prime}$, but optimized towards the original parameters $\theta$. Intuitively, apart from learning the segmentation task on the meta-train data $\mathcal{D}_{s}$, this learning scheme teach the model to learn how to adapt to the domain shift across meta-train data $\mathcal{D}_{s}$ and meta-test data $\mathcal{D}_{t}$.
Consistency Optimization. While the meta-training scheme can simulate domain shifts and help the model adaptation to these shifts, it primarily focuses on making the model robust under shifts within the source domain (as simulated with meta-source and meta-target domains). However, this approach alone is not sufficient to ensure that the model learn the representations that are invariant under different domains. When training a model generalizable for unseen domains, it is crucial to guide the model in learning representations that remain consistent across multiple domains. This is important because, in the feature space, all the labeled data from source domains contribute to training a model that can generalize to previously unseen domains liDomainGeneralizationAdversarial2018 . If the data distributions across various domains within a given feature space remain dissimilar, it hinders the model’s ability to effectively generalize to unseen domains. This is why we propose a consistency optimization method to force the model to learn the domain-invariant representations.
Specifically, consider that a collaborative BEV segmentation model $f(x_{i}^{s};\theta)$ consists of an encoder $E$ and a segmentation decoder head $D_{seg}$, where $x_{i}^{s}$ is the input image and $\theta$ is the model parameters. The encoder $E$ will extract the perception features from different CAVs and fuse the features. The segmentation decoder head $D_{seg}$ will predict the segmentation mask. We formulate the meta-train forward propagation on source domain images as:
$\hat{y}_{i}^{s}=f(x_{i}^{s};\theta)=D_{seg}(E(x_{i}^{s};\theta_{E});\theta_{D})$ (5)
where model parameters $\theta$ are decoupled into $\theta_{E}$ and $\theta_{D}$ for encoder and decoder, respectively. Then, after computing the meta-train loss $\mathcal{L}_{CE}(x_{i}^{s};\theta)$, we can obtain the updated model parameters $\theta^{\prime}$ by the inner loop update (Here, the updated model parameters $\theta^{\prime}$ will just be used in the meta-test step, and will be released after current iteration, the real updated parameters will be computed later). Then we leverage the updated model parameters $\theta^{\prime}$ to compute the meta-test forward propagation prediction on target domain images $x^{t}_{i}$ which are generated by our amplitude augmentation method:
$\hat{y}_{i}^{t}=f(x_{i}^{t};\theta^{\prime})=D_{seg}(E(x_{i}^{t};\theta_{E}^{\prime});\theta_{D}^{\prime})$ (6)
Then, we extract the latent representation from input images $x_{i}^{s}$ and target domain images $x_{i}^{t}$ in meta-train and meta-test step, respectively:
$z^{s}=E(x_{i}^{s};\theta_{E}),~{}z^{t}=E(x_{i}^{t};\theta_{E}^{\prime})$ (7)
Next, in order to enhance the domain-invariant feature capture capability of the model, we design the consistency loss function that leverages the maximum mean discrepancy grettonKernelMethodTwoSampleProblem to make the latent representations $z^{s}$ and $z^{t}$ more similar. We first map the features to a reproducing kernel Hilbert space (RKHS) $\mathcal{F}$, and then compute the distance between their mean in RKHS as follows:
$\mathcal{L}_{cons}=\left\|\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}\phi(z_{i}^{s})-\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\phi(z_{i}^{t})\right\|^{2}_{\mathcal{F}}$ (8)
where $\phi$ is the kernel function that maps the distribution to RKHS, $n_{s}$ and $n_{t}$ are the numbers of samples in source and target domain, respectively. Typically, the Gaussian RBF kernel is used as the kernel function:
$k(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right)$ (9)
Then, the consistency loss function is given by grettonKernelMethodTwoSampleProblem :
$\begin{split}\mathcal{L}_{cons}&=\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}k(z_{i}^{s},z_{j}^{s})+\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}k(z_{i}^{t},z_{j}^{t})\\ &-\frac{2}{n_{s}n_{t}}\sum_{i,j=1}^{n_{s},n_{t}}k(z_{i}^{s},z_{j}^{t})\end{split}$ (10)
Algorithm 1 Meta-consistency Training
1:  Input: Meta-train data $\mathcal{D}_{s}$, Meta-test data $\mathcal{D}_{t}$, Learning rates $\alpha$, $\gamma$, Consistency weight $\beta$
2:  Initialize: Model parameters $\theta$
3:  while not converged do
4:     Meta-Train:
5:     Obtain the prediction using meta-train data $\mathcal{D}_{s}$: $\hat{y}^{s}_{i}=D_{seg}(E(x_{i}^{s};\theta_{E});\theta_{D})$
6:     Obtain the latent feature $z^{s}=E(x_{i}^{s};\theta_{E})$
7:     Compute the cross entropy loss $\mathcal{L}_{CE}$ with prediction $\hat{y}^{s}_{i}$ and ground truth $y_{i}^{s}$
8:     Update parameters $\theta^{\prime}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{CE}(\mathcal{D}_{s};\theta)$
9:     Meta-Test:
10:     Obtain the prediction using meta-test data $\mathcal{D}_{t}$ with updated parameters: $\hat{y}^{t}_{i}=D_{seg}(E(x_{i}^{t};\theta_{E}^{\prime});\theta_{D}^{\prime})$
11:     Obtain the latent feature $z^{t}=E(x_{i}^{t};\theta_{E}^{\prime})$
12:     Compute the cross entropy loss $\mathcal{L}_{CE}$ with prediction $\hat{y}^{t}_{i}$ and ground truth $y_{i}^{t}$
13:     Consistency Optimization:
14:     Compute consistency loss $\mathcal{L}_{cons}$ with two latent features $z^{s}$ and $z^{t}$ by Eq. (10)
15:     Obtain the final updated parameters $\theta\leftarrow\hat{\theta}$ in this iteration by Eq. (12)
16:  end while
Overall learning objective. The overall meta learning objective is composed of the cross entropy loss $\mathcal{L}_{CE}$ and the consistency loss $\mathcal{L}_{cons}$ as follows:
$\mathcal{L}_{meta}=\mathcal{L}_{CE}(x_{i}^{t};\theta^{\prime})+\beta\mathcal{L}_{cons}(x_{i}^{s},x_{i}^{t};\theta^{\prime})$ (11)
where $\beta$ is the balancing weight on the consistency loss, while $\theta^{\prime}$ is the updated model parameters obtained from the inner loop update in Eq. (4). Then, the inner loop objective and the meta learning objective are optimized together with respect to the original parameters $\theta$ as follows:
$\hat{\theta}=\theta-\gamma\nabla\left(\mathcal{L}_{CE}(x_{i}^{s};\theta)+\mathcal{L}_{meta}(x_{i}^{s},x_{i}^{t};\theta^{\prime})\right)$ (12)
The whole training procedure is summarized in Alg. 1.
III-D Intra-system Domain Alignment During Inference
Figure 3: Qualitative comparison of image translation in the RGB and LAB color spaces.
In collaborative perception for CAV driving, we observe that the ego vehicle and other vehicles are situated at different environments, e.g., unbalanced lighting conditions: one vehicle in the shade and the other in the open. Moreover, different types of car cameras can cause chromatic aberration. These phenomena can result in intra-system domain discrepancy, which can degrade the performance of a collaborative perception system, to some extent. In this subsection, we propose a simple yet effective method to align the intra-system domain shift during inference. This method can translate the image style by unifying the pixel distribution in color space. Inspired by reinhardColorTransferImages2001 , we employ image translation based on the LAB color space (LAB color space is another color model) to minimize domain discrepancy. Similar to RGB, LAB color space is also a three-dimensional color space. The three dimensions of LAB color space are $L^{*}$, $a^{*}$ and $b^{*}$. $L^{*}$ represents the brightness of the color, and $a^{*}$ and $b^{*}$ represent the color components. This approach is motivated by two key observations. Firstly, the LAB color space has a broader gamut compared to the RGB color space. Secondly, as demonstrated in Fig. 3, images translated in the LAB color space more resemble the style of target domain images than those translated directly in the RGB color space. Thus, after considering a source domain RGB image $x^{s}$ and a target domain RGB image $x^{t}$, we convert the RGB image to the XYZ color space by:
$\left[\begin{array}[]{c}X\\ Y\\ Z\end{array}\right]=\left[\begin{array}[]{lll}0.4124&0.3575&0.1804\\ 0.2126&0.7151&0.0721\\ 0.0193&0.1191&0.9502\end{array}\right]\left[\begin{array}[]{l}R\\ G\\ B\end{array}\right]$ (13)
Then, we can obtain the LAB color space using the following equation:
$L^{\star}$ $=116f\left(Y/Y_{n}\right)-16$ (14)
$a^{\star}$ $=500\left[f\left(X/X_{n}\right)-f\left(Y/Y_{n}\right)\right]$
$b^{\star}$ $=200\left[f\left(Y/Y_{n}\right)-f\left(Z/Z_{n}\right)\right]$
where $f(t)=t^{1/3}$, if $t>\left(\frac{6}{29}\right)^{3}$, otherwise $f(t)=\frac{1}{3}\left(\frac{29}{6}\right)^{2}t+\frac{4}{29}$, and $X_{n},Y_{n},Z_{n}$ are the coordinate values of the reference white point. The LAB color space is a perceptually uniform color space, which means that a small change in LAB color space corresponds to a small change in the perceived color. Therefore, we can translate the image style by manipulating the LAB color space. After converting the RGB images to the LAB color space, denoted as $x^{s}_{LAB}$ and $x^{t}_{LAB}$, respectively, we first compute the mean and standard deviation of the LAB color channels of $x^{s}_{LAB}$ and $x^{t}_{LAB}$, denoted as $\mu^{s}_{LAB}$, $\sigma^{s}_{LAB}$, $\mu^{t}_{LAB}$, $\sigma^{t}_{LAB}$, respectively. Then, we can obtain the translated image $x^{s}_{LAB}\rightarrow x^{t}_{LAB}$ by:
$\hat{x}^{s}_{LAB}=\sigma^{t}_{LAB}/\sigma^{s}_{LAB}\cdot(x^{s}_{LAB}-\mu^{s}_{LAB})+\mu^{t}_{LAB}$ (15)
After obtaining the translated image $\hat{x}^{s}_{LAB}$, we can convert it back to the RGB color as $\hat{x}^{s}$.
In summary, in ollaborative perception, the ego vehicle first converts its perceived RGB images to LAB color space and obtains $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$. Then, the ego vehicle shares $\mu^{t}_{LAB}$ and $\sigma^{t}_{LAB}$ with other vehicles. Other vehicles can then use these parameters to translate their perceived RGB images to the target domain style. This process helps reduce the domain discrepancy among the collaborative vehicles. After that, other vehicles extract the intermediate features from the translated images and send these features to the ego vehicle. The ego vehicle then fuses these features to predict the final segmentation mask.
Figure 4: Visualization of the foggy dataset. The first row is the original image, the second row is the synthesized foggy image.
Figure 5: Visualization of the fog with different densities.
Figure 6: Visualization of the rainy dataset.
Figure 7: Visualization of the night dataset, the first raw and the third raw is the original images, the second raw and the fourth raw is the synthesized night images.
IV Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed method. Specifically, we choose OPV2V as our fundamental dataset and synthesize four different datasets with different domain discrepancy. We first introduce the datasets and metrics used in our experiments. Then, we compare our method with several state-of-the-art methods. Finally, we conduct ablation studies to evaluate the effectiveness of our method.
IV-A Datasets and Metrics
To assess the domain generalization performance of our proposed method, it is essential to evaluate it across various datasets exhibiting domain discrepancies. However, in the domain of collaborative BEV segmentation, there is no public available datasets suitable for this purpose. Consequently, we have taken the initiative to synthesize four distinct datasets, each characterized with unique domain discrepancies, drawing from the OPV2V dataset xuOPV2VOpenBenchmark2022 as the foundation. These datasets include variations such as foggy-weather, rainy-weather, night conditions, and bright sunny scenarios. The specific attributes and details of these synthesized datasets are elaborated upon below.
OPV2V: In our study, we leverage the OPV2V dataset xuOPV2VOpenBenchmark2022 , a comprehensive dataset curated for integrated perception using V2V communications. This dataset, amassed through the CARLA simulator dosovitskiyCARLAOpenUrban2017 and OpenCDA xuOpenCDAOpenCooperative2021 , comprises 11464 frames of LiDAR point clouds and images, with each frame featuring the minimum of 2 and the maximum of 7 interconnected vehicles. It encompasses 73 distinct scenarios, each averaging a duration of 25 seconds.
Sunny: The initial task in our setting is driving on a sunny day. As the OPV2V dataset is collected at the daytime, so we sample a portion of the OPV2V as our sunny dataset.
Fog: The second task is driving on a foggy day. We synthesize a foggy-weather dataset with different densities of fog (e.g., dense fog and light fog) as shown in Fig. 4 and 5. The foggy weather is generated from atmospheric scattering model narasimhanVisionAtmosphere2008 .
Rain: In addition, we synthesize a rainy-weather dataset with different densities of rain (e.g., heavy rain and light rain) as shown in Fig. 6. Specifically, the dataset is generated by utilizing RainMix, which is a novel data augmentation method. In order to generate the rainy dataset, we first sample a rain map from a public dataset of real rain streaks gargPhotorealisticRenderingRain2006 . We then apply random transformations to the rain map using the RainMix technique. These transformations include rotation, zooming, translation, and shearing, which are randomly sampled and combined. Lastly, the rain maps after transformation are merged with the original source domain images.
Night: Finally, as shown in Fig. 7, we synthesize a night dataset by leveraging a day-to-night generative adversarial network (GAN) zhuUnpairedImagetoImageTranslation2020 ; arrudaCrossDomainCarDetection2019 , which can render the daytime images to night images.
For performance assessment, we utilize the Intersection of Union (IoU) metric to contrast the predicted map with the actual map-view labels.
$IoU=\frac{|BEV_{1}\cap BEV_{0}|}{|BEV_{1}\cup BEV_{0}|}$ (16)
where $BEV_{0}$ represents the ground truth BEV map, $BEV_{1}$ represents the predicted BEV map. The IoU metric is calculated for each class and averaged across all classes to obtain the final IoU score.
TABLE I: Comparison results and evaluation of the domain generalization performance with different domains.
Domain Type
Sunny Fog
AP@IoU(%)
Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022
51.92 36.89 22.96 37.27 40.20 11.54 0.04 18.58
DiscoNet liLearningDistilledCollaboration2021
39.83 53.19 37.37 43.46 33.74 33.68 23.61 30.34
V2VNet wangV2VNetVehicletoVehicleCommunication2020
41.33 51.32 37.21 43.29 35.05 36.06 22.10 31.07
CoBEVT xuCoBEVTCooperativeBird2022
52.16 46.58 38.07 45.60 25.43 22.69 7.16 18.42
Ours
55.05 50.84 39.48 46.96 42.01 41.20 28.33 37.18
Domain Type
Rain Night
AP@IoU(%)
Vehicle Road Lane Average Vehicle Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022
47.93 29.15 0.14 25.74 22.58 9.68 2.22 11.49
DiscoNet liLearningDistilledCollaboration2021
36.83 45.48 29.96 37.42 22.62 5.41 9.42 12.48
V2VNet wangV2VNetVehicletoVehicleCommunication2020
41.38 43.47 29.72 38.19 22.43 14.52 4.73 13.89
CoBEVT xuCoBEVTCooperativeBird2022
50.67 29.33 28.24 36.08 5.53 20.16 2.15 9.28
Ours
49.77 36.52 30.80 39.03 25.78 21.61 12.73 20.04
IV-B Implementation Details
Our model is built on PyTorch and trained on two RTX4090 GPUs utilizing the AdamW optimizer. In the training process, the initial learning rate is set to $2\times 10^{-4}$ and decays by an exponential factor of $1\times 10^{-2}$ with a cosine annealing learning rate scheduler. During meta-consistency training, the meta-step size is set to $1\times 10^{-3}$. The hyperparameter $\alpha$ is empirically set to 0.01 to avoid artifacts on the transformed images. The balance weight $\beta$ of consistency loss is set to 0.1. We set the communication range of CAVs to 70m, and all the CAVs outside the communication range from the ego CAV are ignored. In addition, we first scale the camera images into $512\times 512$ pixels, then we employ CVT zhouCrossviewTransformersRealtime2022 as our backbone to extract the features of each CAV, we employ FuseBEVT xuCoBEVTCooperativeBird2022 to fuse the features from different CAVs, and after that we employ a 6 layer convolutional neural network as the decoder to generate the final BEV segmentation map.
IV-C Comparison Analysis
Experimental setting. In our experiment, we first train the all the models in the original OPV2V dataset for 90 epochs with batch size equal to 1, then we continue to train the model on the synthesized datasets for 20 epochs. During continuing training on the synthesized datasets, we follow the convention in domain generalization literature by employing the leave-one-domain-out strategy. Specifically, we train on $K-1$ distributed source domains and evaluate on the remaining unseen target domain. This approach yields four distinct generalization settings for the collaborative BEV segmentation task.
In assessing the performance of our BEV segmentation rendered by our proposed approach, we compare it with several methods as follows: 1) AttFuse xuOPV2VOpenBenchmark2022 : employs self-attention to fuse the features derived from collaborative vehicles; 2) V2VNet wangV2VNetVehicletoVehicleCommunication2020 : leverages a spatially cognizant graph neural network to consolidate information garnered from all proximate CAVs; 3) DiscoNet liLearningDistilledCollaboration2021 : deploys a teacher-student framework to instruct the model via knowledge distillation; 4) CoBEVT xuCoBEVTCooperativeBird2022 : fuses camera features from a myriad of views and agent data within an underlying transformer framework. For the implementation of these methods, we follow the original papers and use the open code provided by the authors and establish them in the collaborative BEV segmentation setting with the same decoder.
Comparison results. The results of our experiments are summarized in Table I, showcasing the domain generalization performance across different environmental conditions. It is evident from the table that our proposed method outperforms the state-of-the-art techniques across all tested domains, affirming the robustness and efficacy of our approach. In the Sunny domain, our method achieves the highest average precision (AP) of 46.96%, a substantial improvement over the next best performance of 45.60% by CoBEVT. Similarly, in the Fog domain, our method again tops the list with an AP of 37.18%, demonstrating a notable gain over the 31.07% achieved by V2VNet. Moreover, in challenging Rain and Night domains, our approach continues to exhibit superior performance. In the Rain domain, we achieve an AP of 39.03%, outclassing the 38.19% by V2VNet, and in the Night domain, our method significantly surpasses others with an AP of 20.04%, with the next best being 13.89% by V2VNet. The comparative analysis elucidates the benefit of employing a unified domain generalization framework for training and inference in collaborative perception. Our method’s ability to generalize across various domains validates the efficacy of the employed amplitude augmentation and meta-consistency training scheme. The results also underline the importance of intra-system domain alignment during the inference phase, significantly reducing the domain discrepancy among CAVs and consequently enhancing the collaborative perception performance across different environmental conditions.
Figure 8: Visualization of the BEV segmentation results of different methods. In the first row, The four images are from ego vehicle’s cameras that in four different angles. In the second row, the five BEV maps are ground truth and generated from different method: column (a) is the ground truth, (b) is generated from AttFuse method, (c) is from DiscoNet, (d) is from V2VNet, (e) is predicted by our method. The third row and forth row are another scene. Compared with other methods, our method shows robust performance in different scenes.
Qualitative analysis. In order to provide a qualitative comparison across different methods, we visualize some BEV segmentation results of different methods (including AttFuse, DiscoNet, V2VNet) in Fig. 8. As we can see in the BEV prediction maps, our model yields perception results that stand out in terms of both comprehensiveness and accuracy when compared with other methods. In both scenes, the AttFuse method exhibits significant omissions in road surface and lane. DiscoNet and V2VNet also more or less fail to segment the complete road and lane. As for the vehicle class, these three compared methods occasionally miss segments and display ambiguous boundaries. In contrast, as seen in Fig. 8 (e), our method can segment the road, lane and vehicle class more accurately and comprehensively. These results can show the superiority of our method in the collaborative perception task.
IV-D Ablation Study
Figure 9: Effect of the meta-consistency training ”w/” means with meta-consistency training, ”w/o” means without meta-consistency training.
Figure 10: Visualization of the consistency loss $\mathcal{L}_{cons}$, meta-test loss and meta-train loss.
Effect of amplitude augmentation. In order to evaluate the effectiveness of the proposed amplitude augmentation, we set the sunny dataset as the unseen domain, and test its performance. Then we conduct an experiment depicted in Table II. We observe that by simply removing the amplitude augmentation, the performance degrade. The IoU accuracy of vehicles, road, lane are decreased by 9.51%, 5.58%, 3.51%, respectively. To further indicate the effect of the AmpAug mechanism and show that AmpAug can increase the diversity of the source domain. We use t-SNE111t-SNE: t-distributed stochastic neighbor embedding is a statistical method for visualizing high-dimensional data by giving each datapoint a location in a two or three-dimensional map. van2008visualizing to visualize the distribution of the original images and the generated images in Fig. 11. Specifically, the pink points denote the local data, the other blue points denote the transformed data that are generated with amplitude spectrum from our constructed dataset $\mathcal{D}_{t}$ proposed in Sec. III-B. From the observation, we can see that the distribution of the transformed images is more diverse than the original images. This indicates that the AmpAug mechanism can increase the diversity of the source domain by augmenting the low-level distribution of the images, thereby making the model to see more low-level variations. This can prevent the model from overfitting to the source domain and help the model to learn the domain-invariant features.
Figure 11: Visualization of the t-SNE embedding for the original dataset (pink points) and the corresponding transformed images by amplitude augmentation dataset (blue points).
TABLE II: Ablation Study Results of the Amplitude Augmentation. “w/” means the with amplitude augmentation, “w/o” means the without amplitude augmentation.
AP@IoU(%) Vehicle Road Lane Average
w/o AmpAug 45.54 45.26 35.97 42.26
w/ AmpAug 55.05 50.84 39.48 48.46
TABLE III: Ablation Study Results of the Intra-System Domain Alignment. “w/” means the with intra-system domain alignment, “w/o” means the without intra-system domain alignment.
AP@IoU (%) Vehicles Road Lane Average
AttFuse xuOPV2VOpenBenchmark2022 w/o 51.92 36.89 22.96 37.27
w/ 53.16 41.09 25.60 39.95
DiscoNet liLearningDistilledCollaboration2021 w/o 39.83 53.19 37.37 43.46
w/ 41.94 53.23 37.80 44.32
V2VNet wangV2VNetVehicletoVehicleCommunication2020 w/o 41.33 51.32 37.21 43.29
w/ 44.85 53.49 39.13 45.82
Ours w/o 54.09 49.13 37.66 46.96
w/ 55.05 50.84 39.48 48.46
Effect of meta-consistency training. In order to evaluate the effectiveness of the proposed meta-consistency training, we conduct an experiment depicted in Fig. 9. We observe that the meta-consistency training does improve the performance of the model: the IoU accuracy of vehicles, road, lane are improved by 2.56%, 1.08% and 0.26%, respectively. Without our method, the model could not learn the domain-invariant features, which leads to the performance degradation. Notably, the performance of vehicle class is more sensitive to the domain discrepancy than the other classes, because vehicles are more dynamic than the other classes. The results demonstrate that our meta-consistency training can improve the performance of collaborative perception and make the model more robust to the domain shift.
To further analyze the effectiveness of the meta-consistency training, we visualize the consistency loss, denoted as $\mathcal{L}_{cons}$ in Fig. 10. It can be observed that despite the consistency loss $\mathcal{L}_{cons}$ has some fluctuations, it shows the decreasing trend during the training process. This suggests that the consistency loss can make the latent representations more similar, thereby aiding the model to learn the domain-invariant features. In addition, the meta-test loss and meta-train loss also show a decreasing trend. The meta-test loss has a large fluctuation at the beginning during the training, but it gradually levels off around 0.06.
Effect of intra-system domain alignment. In order to evaluate the effect and the university of our intra-system domain alignment mechanism, we conduct an experiment with different collaborative perception framework, including AttFuse xuOPV2VOpenBenchmark2022 , V2VNet wangV2VNetVehicletoVehicleCommunication2020 , and DiscoNet liLearningDistilledCollaboration2021 . Before evaluating on these different collaborative perception frameworks, we first implement our method on each framework. The results are shown in Table III. We can see that our intra-system domain alignment mechanism can generally improve the performance of the baselines, especially for the vehicle class. Specifically, for the AttFuse method, the IoU accuracy of vehicles, road and lane are improved by 1.24%, 4.23%, 2.64%, respectively. For the V2VNet, the IoU accuracy of vehicles, road and lane are improved by 3.52%, 2.17%, 1.92%, respectively. In addition, as for our method, we simply remove this component and evaluate, we can see that the performance of the model will be degraded. These results can demonstrate the effect of our intra-system domain alignment mechanism that can generally improve the performance of the collaborative perception framework by reducing the distribution heterogeneity among the data in different vehicles.
V Conclusion
In this paper, we have developed a novel framework to tackle the domain generalization with collaborative perception systems for connected and autonomous driving. We first propose an amplitude augmentation method to augment the original source domain data to different target domains in collaboration with our constructed dataset. Then, we have designed a meta-consistency training scheme based on meta-learning paradigm. This training scheme can simulate the domain shift, guiding the model to learn how to learn from different domains, and the meta-consistency loss can force the model to learn the domain-invariant features, thereby enhancing the ability of generalization of the model. Finally, we have leveraged the intra-system domain alignment method to minimize the domain gap among the data perceived by different collaborative vehicles prior to inference. Comprehensive experiments have demonstrated the superiority of our framework compared with the existing state-of-the-art methods.